### v1

@@ 3.1 Methodologies
There are multiple methodological approaches used in human-AI interaction research. These methodologies can broadly be categorized into qualitative and quantitative approaches. Quantitative methods, in turn, can be further divided into survey-based studies and experimental research designs. Among 89 papers reviewed in this study, 12 employed a survey methodology, 31 followed an experimental design, and 38 utilized qualitative methods.
@@
Additionally, 20 of the reviewed studies did not fit neatly into the qualitative-quantitative dichotomy. Examples include case studies in which the authors both describe their own experiences and provide quantitative metrics on system behavior (Ross et al., 2023b), quantitative analysis of system behavior using pre-defined datasets of typical user tasks (Nguyen and Nadi, 2022; Fagadau et al., 2024), and investigations of public discussions on platforms such as Reddit (Klemmer et al., 2024) and GitHub Issues (Zhou et al., 2025).
@@
The sample sizes for different methodological approaches varied considerably. For interview-based studies, the sample size ranged from 7 to 61 participants (median = 16). Experimental designs involved sample sizes ranging from 17 to 214 participants, with the exception of two A/B studies, one with 535 participants (Mozannar et al., 2023) and another, conducted by Meta (Dunay et al., 2024), reporting “thousands” of participants. As expected, survey-based studies had the largest sample sizes, ranging from 68 to 2,047 participants (median = 507).
3.2 Context of studies
The studies analyzed in this review represent two primary contexts for working with AI in software development: professional (82%) and educational (18%). These contexts provide contrasting perspectives on the integration of AI tools, with professional studies focusing on real-world applications and educational studies exploring their role in learning environments.
The majority of professional studies emphasize the practical utility of AI tools. Most studies address coding (26%) and maintenance (9%), with limited exploration of requirements gathering (1%) or other SDLC stages. However, 45 studies (61%) do not specify the SDLC stage, indicating a gap in contextual precision.
Educational studies primarily focus on pedagogical insights. These studies often investigate how students use AI tools to learn programming or solve computational problems. While a few studies address coding and requirements gathering (15%), most (81%) do not specify the SDLC stage, aligning with the broader exploratory nature of educational research.
This distribution reveals notable gaps in the literature:
  • Lack of perspectives from non-adopters of AI tools for development.
  • Limited number of educational studies highlighting the need for a deeper exploration of AI ’s impact on CS education, especially as developers constantly learn and re-learn programming practices.
A subset of studies (5%) conceptualizes programming with AI as a form of pair programming, where AI acts as a collaborative partner. This perspective highlights the evolving view of AI as not just a tool but a collaborator, representing a growing trend in co-creative programming workflows and suggesting new opportunities and challenges for both practitioners and educators.
3.3 Design
One of the research areas that emerged from both our current review and our previous work (Sergeyuk et al., 2024) is the Design of HAX in IDE. Among the reviewed studies, approximately 31% investigate forms of AI integration into IDEs, with most of them (22 out of 28) explicitly examining how these forms of integration impact users.
The research presents two primary paradigms of AI-assisted interaction inside IDEs: autocompletion-based assistance and conversational agents. Autocompletion interfaces provide short, low-friction code completions, enabling developers to work seamlessly with inline suggestions. These approaches minimize cognitive load and support rapid decision-making without disrupting workflow (Vaithilingam et al., 2023; Yen et al., 2023). In contrast, conversational interactions — often embodied in chat-based assistants or virtual avatars — facilitate higher-level reasoning, iterative refinement, and complex problem-solving(Ross et al., 2023a; Robe and Kuttal, 2022; Gu et al., 2023). Rather than being opposing paradigms, these interaction modes are complementary (Weber et al., 2024), with studies advocating for hybrid models that seamlessly integrate autocompletion with interactive, context-sensitive conversations. In addition, some studies experiment with alternative modalities of interaction that go beyond these two dominant paradigms. For example, mixed-reality interfaces(Manfredi et al., 2023) introduce conversational AI in augmented pair programming environments, leveraging immersive learning experiences for skill development.
Due to the conversational nature of interaction with AI in IDE, prompting became a prominent object of investigations with 12% of studies focusing on this aspect. Across reviewed studies, prompt engineering is positioned as a strategic skill that shapes how developers and AI systems collaborate, affecting both individual productivity and team-based workflows. A central challenge in AI-assisted development is ensuring that AI-generated code aligns with user intent. Studies investigating prompt formulation (Fagadau et al., 2024; Denny et al., 2023) demonstrate that structured and well-contextualized prompts lead to more accurate and reliable output, while excessive detail can degrade performance. As AI tools become embedded in professional and educational workflows, collaborative prompt engineering is also gaining importance. Some works (Feng et al., 2024; Cheng et al., 2024b) introduce frameworks for shared prompt refinement, knowledge transfer, and workflow optimization, suggesting the need for modularity and re-usability in AI interactions.
Beyond interaction paradigms, around 10% of the studies explore design principles for AI-powered IDEs. These principles guide how AI should be embedded into developer workflows while maximizing usability, reliability, and efficiency:
  • Context Awareness: LLM-powered assistance significantly boosts task completion rates when it can effectively incorporate a developer’s code as context (Nam et al., 2023b).
  • Explainability & Transparency: Tools should provide context-aware explanations and highlight uncertain AI-generated code (Vasconcelos et al., 2024; Yan et al., 2024; Sun et al., 2022; Wang et al., 2024) to improve user trust.
Among the specific implementations of AI for IDEs, GitHub Copilot is by far the most studied, with  53% of works investigating it. This concentrated attention on a single tool suggests a risk of overgeneralization, where findings may not fully translate to other AI-powered coding assistants, limiting broader design and implementation insights.
3.4 Impact
Another notable and prominent branch of research is dedicated to Impact of HAX in IDE with 82% of reviewed studies focusing on it.
The increasing integration of AI-powered tools in IDEs has reshaped the way developers engage with coding, debugging, documentation, and software design. AI ’s role extends beyond simple automation, influencing productivity, cognitive processes, and software engineering practices. Although AI has undoubtedly enhanced productivity in some contexts, its overall effects remain highly context-dependent and nuanced. Studies consistently highlight the dual nature of AI’s impact, providing both benefits and challenges.
Multiple studies have confirmed a notable boost in productivity when developers incorporate AI into their workflows. Potential benefits include a higher “ceiling” of productivity for experienced developers and faster onboarding for newcomers (Vasiliniuc and Groza, 2023). For instance, participants using GitHub Copilot completed their tasks up to 55.8% faster than control groups (Peng et al., 2023), while other investigators observed productivity gains ranging from 26% to 35% (Pandey et al., 2024). These improvements are partially attributed to fewer context switches (less time navigating external resources) and reduced boilerplate coding, enabling developers to offload repetitive tasks to AI while focusing on higher-level logic. However, gains are inconsistent: tasks involving proprietary or highly complex logic see only marginal benefits, as AI struggles with context-specific nuances.
Although AI-driven suggestions frequently accelerate development, they also require new verification efforts. When using AI tools, verifying suggestions, refining prompts, and reworking code generated by AI might take up to 50% of developers’ time (Mozannar et al., 2024a). This increased mental overhead arises because AI ’s output can be partially correct yet subtly flawed (Wermelinger, 2023), requiring careful review and, at times, stepwise interaction or re-prompting (Kazemitabaar et al., 2024). Such extra validation often mitigates the risk of blindly accepting erroneous suggestions — a phenomenon labeled “automation bias” (Al Madi, 2022).
Studies focusing on novice programmers reveal both benefits and pronounced risks (Prather et al., 2023; Puryear and Sprint, 2022). Beginners often solve tasks quicker with AI’s help. However, they can over-trust AI, losing sight of important foundational concepts (Prather et al., 2024). In certain cases, when system suggestions are misleading, novices drift away from correct solutions, as they lack the expertise to detect subtle errors (Zhou and Li, 2023). However, structured prompts and explanatory interfaces can help novices understand AI outputs better, fostering safer learning pathways (Kruse et al., 2024).
While most studies focus on short-term task performance, there is an emerging interest in longitudinal effects. Areas of concern include skill atrophy (if developers delegate too much to AI), shifts in collaboration norms among team members, and changes in programming practices.
3.4.1 Attitude
Some studies focus on the impact of HAX in the form of user attitudes toward AI systems integrated in IDEs.
Trust in AI-powered tools is heavily influenced by both the characteristics of AI outputs and user expertise (Amoozadeh et al., 2024; Brown et al., 2024). High-quality, contextually relevant suggestions improve trust, while issues such as inconsistent or overly complex recommendations hinder it. Other situational factors, such as task complexity or file type, also shape developer trust. Tools that fail to support developers in assessing the trustworthiness of outputs often lead to cautious or dismissive behaviors, while explainability layers, usage statistics, and control mechanisms have been proposed to enhance trust-building (Wang et al., 2024).
An important challenge related to trust in AI tools is over-reliance on AI-generated output. Over-reliance occurs when developers accept AI-generated outputs without sufficient validation, posing potential risks. Studies indicate that users tend to overestimate the reliability of AI-generated code (Al Madi, 2022). Conversely, some studies highlight under-reliance, where developers reject AI outputs due to complexity or lack of explainability, even when the code is correct (Sun et al., 2022). Live programming environments, which enable real-time validation, have shown potential in mitigating both over-reliance and under-reliance by lowering the cognitive effort needed for review (Ferdowsi et al., 2024).
3.5 Quality
Ensuring the Quality of AI-assisted code generation is a central concern of 21% of the papers in the reviewed literature. Across studies, quality is examined from multiple perspectives, including code correctness, security, and readability.
One recurring theme is the trade-off between efficiency and correctness. AI-assisted coding significantly accelerates development workflows, but at the cost of increased susceptibility to subtle errors, security vulnerabilities, and reduced maintainability (Sandoval et al., 2023; Tang et al., 2024; Pearce et al., 2021). Several studies suggest that while AI-generated code often appears syntactically valid, it can contain logical flaws, insecure patterns, or deviations from best practices (Pearce et al., 2021; Asare et al., 2024). As a result, effective quality control mechanisms, such as automated verification, static analysis, and model-aware debugging tools, are emphasized as critical to responsible adoption (Agarwal et al., 2024).
Beyond correctness, readability and maintainability emerge as key aspects of AI-generated code quality. Although AI can generate complex solutions rapidly, its output is not always optimized for human understanding. Studies highlight that AI-generated code may be less readable due to overly concise structures, unconventional variable naming, or lack of meaningful comments (Al Madi, 2022). This raises concerns about long-term maintainability, particularly in team-based software development, where code legibility is as important as functional correctness. Research suggests that integrating AI-powered refactoring and explanation tools can help mitigate these challenges, enhancing both short-term usability and long-term code quality (Yan et al., 2024).
Security poses an additional concern: up to 36% of vulnerabilities in AI-assisted code originate from the models themselves, pointing out the risks of replicating insecure patterns from training data (Sandoval et al., 2023). However, programmers remain a major source of vulnerabilities, emphasizing the need for cautious oversight of AI suggestions.
Finally, quality in AI-assisted development is not just a technical issue but also a human and organizational challenge. Research highlights that quality outcomes are influenced by how developers interact with AI tools: whether they verify AI-generated code, how they prompt AI models, and whether they develop strategies for AI-assisted debugging (Weber et al., 2024; Mozannar et al., 2024c; Sandoval et al., 2023; Al Madi, 2022; de Moor et al., 2024). Best practices for improving AI-assisted development quality extend beyond tool design and include education, training, and workflow adaptation.
While LLM-powered tools improve software development, maintaining consistently high-quality code remains a challenge. Ensuring reliability requires better alignment with developers’ needs and robust verification mechanisms (Tang et al., 2024; Agarwal et al., 2024). Studies highlight the need for strategies such as filtering suboptimal suggestions, improving contextual prompting through docstrings and function names, and leveraging adaptive personalization techniques (Mozannar et al., 2024c; de Moor et al., 2024).
Future research calls for more diverse benchmarks, longitudinal studies assessing real-world use, and broader quality evaluations that account for maintainability, reliability, and security (Sandoval et al., 2023; Pearce et al., 2021). Advancing AI-assisted programming will depend on integrating these refinements to ensure not only efficiency, but also long-term software quality.
3.6 Future work proposed by the studies
By analyzing the future work proposed by each study alongside the goals of others, we identified alignment patterns and gaps that highlight both the progress made and the opportunities missed.
@@ 3.6 Future work proposed by the studies
Some future work suggestions align with the goals of other studies, but these connections are not always explicit, reflecting a fragmented research landscape. Common themes like productivity, security, and usability appear across studies but are addressed in isolated contexts without direct cross-referencing. This lack of continuity limits the field’s ability to build cumulative knowledge effectively.
A significant portion of the suggested directions for future research remains unaddressed, indicating how young the field is and highlighting a variety of opportunities for the academic community. Key gaps persist, particularly in studying personalization opportunities, adoption barriers, and cross-context comparisons. Moreover, many studies suggest conducting longitudinal research or involving larger sample sizes. Furthermore, ethical challenges such as bias mitigation, misuse prevention, and AI governance in software development remain largely unresolved.
Addressing these gaps through deliberate cumulative and interdisciplinary research will not only advance the field but also ensure its continued relevance in shaping AI-enhanced software development.

### v2

@@ 3.5 Methodologies
There are multiple methodological approaches used in human-AI interaction research. These methodologies can broadly be categorized into qualitative and quantitative approaches. Quantitative methods, in turn, can be further divided into survey-based studies and experimental research designs. Among 90 papers reviewed in this study, 12 employed a survey methodology, 32 followed an experimental design, and 38 utilized qualitative methods.
@@
Additionally, 20 of the reviewed studies did not fit neatly into the qualitative-quantitative dichotomy. Examples include case studies in which the authors both describe their own experiences and provide quantitative metrics on system behavior (Ross et al., 2023b), quantitative analysis of system behavior using pre-defined datasets of typical tasks (Nguyen and Nadi, 2022; Fagadau et al., 2024), and investigations of public discussions on platforms such as Reddit (Klemmer et al., 2024) and GitHub Issues (Zhou et al., 2025b).
@@
The sample sizes for different methodological approaches varied considerably. For interview-based studies, the sample size ranged from 7 to 61 participants (median = 16). Experimental designs involved sample sizes ranging from 17 to 214 participants, with the exception of two A/B studies, one with 535 participants (Mozannar et al., 2024b) and another, conducted by Meta (Dunay et al., 2024), reporting “thousands” of participants. As expected, survey-based studies had the largest sample sizes, ranging from 68 to 2,047 participants (median = 507).
3.6 Future Work Suggested by Existing Literature
To address RQ3, we analyzed 250 future work statements extracted from the reviewed papers (see the full table in (Sergeyuk et al., 2025)). Each item was open-coded by topic (what is proposed) and method (how it is to be studied), resulting in 17 topics and 10 methodological strategies.
Figure 6: Paper IDs for Topics of suggested future work.
In Figure 6, we present corresponding Paper IDs for each topic, and Figure 7 visualizes the distribution of papers across combinations between topics and methods.
Figure 7: Paper IDs for co-occurrence between Topic on the y-axis and Method on the x-axis for suggested future work.
@@ 3.6 Future Work Suggested by Existing Literature
By analyzing the future work proposed by each study alongside the goals of others, we identified alignment patterns and gaps that highlight both the progress made and the opportunities missed. Some future work suggestions align with the goals of other studies, but these connections are not always explicit, reflecting a fragmented research landscape. Common topics like productivity, audit of AI-generated code, and usability of designed AI tools appear across studies but are addressed in isolated contexts without direct cross-referencing. This lack of continuity limits the field’s ability to build cumulative knowledge effectively.
Suggestions cluster in a set of 17 themes. The most common address productivity factors (43), the design of AI assistance (29), and audits of AI-generated code (28). Additional areas include prompting support (19), skill building with AI (17), education with AI (16), personalization (16), explainability and transparency (15), verification support (13), and broader coverage of the SDLC (12). Less frequent but present are trust and reliability (11), IDE redesign for AI (10), assistant context enrichment (6), user mental models (5), proactivity (4), user control (3), and AI governance (3).
Productivity-oriented suggestions emphasize broader sampling and comparative designs, often at a larger scale (Al Madi, 2022; Prather et al., 2023; Angert et al., 2023; Jiang and Coblenz, 2024; Rasnayaka et al., 2024). They ask for studies with professional developers in realistic settings (Peng et al., 2023; Li et al., 2024; Jiang and Coblenz, 2024; Rasnayaka et al., 2024; Zhu et al., 2024; Kazemitabaar et al., 2025) and for attention to students and novices (Jiang and Coblenz, 2024; Amoozadeh et al., 2024; Kazemitabaar et al., 2025). Authors also recommend evaluations that account for project and environment factors Peng et al. (2023); Feldman and Anderson (2024); Zhou et al. (2025b). Several papers connect productivity to interface and workflow choices and propose concrete design changes and controls (Ziegler et al., 2022; Angert et al., 2023; Rasnayaka et al., 2024; Ságodi et al., 2024; Pandey et al., 2024; Amoozadeh et al., 2024; Nam et al., 2023; Jiang and Coblenz, 2024).
Future work on auditing AI-generated code seeks wider security coverage, longitudinal observation, and stronger assets for evaluation. Proposals include expanding vulnerability categories and language coverage and tracking change over time (Pearce et al., 2022; Asare et al., 2023, 2024), building targeted test suites and benchmarks (Liu et al., 2024), and increasing unit tests and quality metrics to stress code quality (Yetistiren et al., 2022; Liu et al., 2024). Some papers extend audits to SDLC touchpoints and to interfaces that surface risk at decision time (Liang et al., 2024; Asare et al., 2024; Liu et al., 2024; Kruse et al., 2024; Pearce et al., 2022; Vasconcelos et al., 2025).
Prompting support centers on guided interactions and safe practice. Suggestions include structured conversational prompting and refinement aids (Jiang et al., 2022; Angert et al., 2023), prompt engineering strategies with comparative evaluation (Pearce et al., 2022; Denny et al., 2023; Mastropaolo et al., 2023; Fagadau et al., 2024), multi-modal prompt channels for tasks that are not purely textual (Angert et al., 2023), and prompt sanitization with privacy-aware feedback (Li et al., 2024). Authors also ask for generalization across tools and settings (Fagadau et al., 2024).
Personalization proposals focus on adapting assistance to user profiles, expertise, and task state. Examples include user models to personalize initial prompts and responses (Ross et al., 2023a; Chopra et al., 2024), frequency and snooze controls to manage attention (Vaithilingam et al., 2023), adaptive strategies that respond to real time performance and needs (Kazemitabaar et al., 2025), and results that better match the intended coding context and style (de Moor et al., 2024; Liu et al., 2024; Sahoo et al., 2024; Cheng et al., 2024b; Mozannar et al., 2024c; Vasconcelos et al., 2025). Several papers request tailored documentation and scaffolds for specific roles such as data analysts (Wang et al., 2022; Gu et al., 2024).
Explainability and transparency aim to help users understand, validate, and teach the assistant. Suggested mechanisms include rationales, annotations, and code level explanations (Vaithilingam et al., 2022; Hu et al., 2022; McNutt et al., 2023), interpretability features that support accurate mental models and calibrated expectations (Weisz et al., 2022; Jiang et al., 2022; Sun et al., 2022; Bird et al., 2022), internal deliberation to improve reasoning quality (Ross et al., 2023b), and presentation of validation evidence such as testing suite results or multi model checks (Omidvar Tehrani et al., 2024; Yan et al., 2024).
Verification support concentrates on tools that help users check and repair outputs before adoption. Proposals include automatic test generation and retrieval of reference examples (Vaithilingam et al., 2022; Ferdowsi et al., 2024), post processing and filtering layers that repair or block risky outputs (Pearce et al., 2022; Al Madi, 2022; Ferdowsi et al., 2024), interface patterns that present multiple alternatives for comparison (Weisz et al., 2022), and study designs that measure how learners verify AI suggestions and develop durable checking habits (Kazemitabaar et al., 2023b, a).
Coverage beyond the implementation stage remains a consistent request. Papers propose research and tooling for requirements, design, evolution, debugging, testing, deployment, and code search across the life cycle (Nguyen and Nadi, 2022; Liu et al., 2024; OBrien et al., 2024; Chopra et al., 2024; Ferdowsi et al., 2024; Vasconcelos et al., 2025). Many of these suggestions pair life cycle scope with deeper integration into the developer workspace (Chopra et al., 2024; Ferdowsi et al., 2024).
Trust and reliability suggestions examine individual and community factors as well as interface effects. Authors ask for studies of how communities and organizations form and recalibrate trust over time (Cheng et al., 2024a; Prather et al., 2024), for quantification of interface-driven trust shifts (Wang et al., 2024), for evaluations that measure the impact of assistant behavior changes on acceptance (Feldman and Anderson, 2024), and for approaches that foster appropriate trust with clear boundaries (Ferdowsi et al., 2024). Inclusion of diverse developer groups appears throughout (Vaithilingam et al., 2022; Hu et al., 2022; Ross et al., 2023a).
Proposals on IDE redesign and context enrichment argue for AI-aware environments. Suggestions include non-linear input and new interaction media (McNutt et al., 2023), improved error handling and better merging of sketch-like artifacts (Angert et al., 2023), chat or dialogue interfaces for automation control (Shlomov et al., 2024), and consistent propagation of edits and assumptions (Kazemitabaar et al., 2024a). For context, papers recommend memory management, dynamic prompts that reflect project artifacts, and search-based integrations (Ross et al., 2023b; Wang et al., 2023b; Ross et al., 2023a; Shlomov et al., 2024).
Less frequent but important themes include user mental models, proactivity, user control, and governance. Future work asks for theory building on how developers form and use mental models of assistants and for longitudinal observation of expectation change (Nguyen and Nadi, 2022; Lau and Guo, 2023; Asare et al., 2023; Zhu et al., 2024; Amoozadeh et al., 2024). Proactivity-related suggestions seek predictive models of user state and careful study of long-term effects on quality of code and productivity (de Moor et al., 2024; Mozannar et al., 2024a, b). Use control-oriented suggestions call for clearer and richer controls over generation and automation levels (Yen et al., 2023; Feng et al., 2024; Wang et al., 2024). Governance suggestions raise questions about fairness, access, and alignment, including alignment toward helpful and harmless behavior and possible career impacts for different demographic groups (Ross et al., 2023a; Feldman and Anderson, 2024; Peng et al., 2023).
Taken together, these directions indicate the need for larger and longer evaluations, stronger audit and verification assets, broader life cycle coverage, and adaptive assistance that is explainable and under user control.